OpenStudy (idealist10):

Solve for y: y+y^2=x^2+c

OpenStudy (idealist10):

@amistre64

OpenStudy (amistre64):

you have a quadratic in y

OpenStudy (amistre64):

y^2 + y - (x^2 + c) = 0

OpenStudy (idealist10):

And then?

OpenStudy (amistre64):

and then use the quadratic formula ...

OpenStudy (idealist10):

Let me work it out.

OpenStudy (amistre64):

ill let you :)

OpenStudy (idealist10):

I got y=-1+/-sqrt(4x^2+4c+1) divided by 2 but if I plug in 2 into x, I got c=-4. The answer for this problem should be y=-1+sqrt(4x^2-15) divided by 2.

OpenStudy (idealist10):

It's an initial-value problem where y(2)=0.

OpenStudy (amistre64):

\[y^2 + y - (x^2 + c) = 0\] \[y=\frac{-1\pm\sqrt{1^2+4(x^2+c)}}{2}\] \[y=\frac{-1\pm\sqrt{4x^2+c}}{2}\]

OpenStudy (amistre64):

im assuming c is an all encompassing constant

OpenStudy (idealist10):

But why doesn't the answer doesn't contain +/- and only + sign?

OpenStudy (amistre64):

\[0=\frac{-1\pm\sqrt{16+c}}{2}\] \[0=-1\pm\sqrt{16+c}\] \[1=\sqrt{16+c}\] c = -15

OpenStudy (amistre64):

its just a matter of the mathing process. 1 is a postive number, and there is no sqrt value that is negative

OpenStudy (amistre64):

no 'real' sqrt ...

OpenStudy (idealist10):

So what's the correct answer? The one that doesn't contain +/- sign and only the plus sign?

OpenStudy (amistre64):

the +- sign is an operation that allows us to + or - the sqrt as needed. it just so happens that in this case there is no subtraction operation that produces a valid real solution

OpenStudy (amistre64):

\[1 = \pm\sqrt{16+c}\] \[1=\sqrt{16+c}~~or~~1=-\sqrt{16+c}\] since there is not sqrt(n) that gives us a negative value, the right option is simply not valid

OpenStudy (idealist10):

Thanks!

OpenStudy (amistre64):

yw

OpenStudy (idealist10):

:)